can you please help me :
convert this linear equation to standard form?

y=1/2x-2/3

Answer by ankor@dixie-net.com(22740)

y = 1/2x - 2/3
:
Get rid of the fractions, multiply both sides by 6
6y = 3x - 4
:
Subtract 3x from both sides, results
-3x + 6y = -4
Multiply equation by -1 to change the signs
3x - 6y = 4